OpenStudy (anonymous):

what are the possible values for sin theta in: 2sin^2 theta + sin theta -1 = 0

OpenStudy (shamim):

\[2 \sin ^{2} \theta +\sin \theta -1=0\]

OpenStudy (shamim):

\[2 \sin ^{2} \theta + 2 \sin \theta - \sin \theta-1=0\]

OpenStudy (shamim):

\[2 \sin \theta (\sin \theta +1)-(\sin \theta +1)=0\]

OpenStudy (shamim):

\[(2 \sin \theta-1)(\sin \theta +1)=0\]

OpenStudy (shamim):

please do the rest

OpenStudy (shamim):

if u face any problem plz ask me

OpenStudy (anonymous):

cool! thanks

OpenStudy (shamim):

welcome